OpenStudy (anonymous):

Please help!!!! A cat jumps straight up from the ground with an initial vertical velocity of 9 feet per second. a. Write an equation that models the height [h] (in feet) of the cat as a function of the time [t] (in seconds) since it jumps. b. After how many seconds does the cat land on the ground?

OpenStudy (anonymous):

could have sworn i answered this already. was it wrong?

OpenStudy (anonymous):

No, but I didn't understand part B. :(

OpenStudy (amistre64):

v(0) = -32(0) + C = 9 ; C = 9 v(t) = -32t +9 h(t) = -16t^2 +9t + C; where h(0) = 0, C = 0 h(t) = -16^2 +9t

OpenStudy (amistre64):

hah ... loped off a t

OpenStudy (anonymous):

c be 0

OpenStudy (amistre64):

h(t) = -16t^2 +9t now the height of the cat is given at any moment in time; so when it lands what height is it gonna be at?

myininaya (myininaya):

is the height 0 when it lands?

OpenStudy (amistre64):

if 0 height is on the ground; at what height do we need to determine for the cat to be back on the ground?

OpenStudy (anonymous):

i think the confusion was how to solve this we wrote \[-16t^2+9t=0\] right? factor as \[t(-16t+9)=0\] then \[t=0\] or \[-16t+9=0\] \[9=16t\] \[\frac{9}{16}=t\]

OpenStudy (amistre64):

it takes 9 years and 16 weeks to land again ;)

OpenStudy (amistre64):

it helps, since we are able to have real interactions between questioner and answerers, that if you dont understand something that all you have to do is just ..... well, say something. reposting a question simply becasue you dont understand it doesnt really help out in that regards. It just makes those that took their free time to try to help you out feel rather unappreciated

OpenStudy (amistre64):

just dont be afraid to have a real conversation, if that is the case :)

OpenStudy (anonymous):

Sorry :(